James walks a distance of 10 km at an average speed of x km/h. Jane walks the same distance at (x - 1) km/h. If the difference in their times is 30 minutes, find the value of x.
Answer by mananth(16946)


James walks a distance of 10 km at an average speed of x km/h. Jane walks the same distance at (x - 1) km/h. If the difference in their times is 30 minutes, find the value of x.
James walks a distance of 10 km at an average speed of x km/h
time taken by James = 10/x
Jane walks the same distance at (x - 1) km/h
Time taken by Jane = 10/(x-1)
Difference in time = 30 minutes = 1/2 hour
10/(x-1)-10/x = 1/2
Multiply by the LCD 2x(x-1)
10.2(x)-10*2*x = x(x-1)
20x-20x -20= x^2-x
x^2-x +20=0
x^2-5x-4x+20=0
x(x-5)-4(x-5)=0
(x-5)(x-4)=0
x= 5 or 4
4 not possible
x=5